# Distal Causal Excursion Effects: Modeling Long-Term Effects of Time-Varying Treatments in Micro-Randomized Trials

### v1

@@ A.4 Example 4 (extension): DCEE captures treatment effects on future eligibility
  • $-p\rho_{0}\alpha$: the negative interaction between $A_{1}$ and $A_{2}$ not accounting for the influence of $A_{1}$ on $I_{2}$;
@@
  • $-p\rho_{1}\beta_{2}$: the reduced effectiveness of $A_{2}$ on $Y$ due to $A_{1}$ negatively impacting future eligiblity $I_{2}$.
Appendix B Proof of Identification Result (Theorem 1)
@@
We now prove Theorem 1.
Proof of Theorem 1.
@@
Here, Eq. 13 follows from Lemma 3, Eq. 14 follows from Assumption 1 (i), Eq. 15 follows from Assumption 1 (iii), Eq. 16 follows from Assumption 1 (iii), Eq. 17 follows from Assumption 1 (i).
The fact that Eqs. 12, 13 and 17 are equal to each other establishes Eqs. 3 and 4.
@@ Appendix C The Weighted Average Form of $\beta^{\star}$
×$\times\sum_{t=1}^{T}\omega(t)\mathbb{E}(I_{t})\mathbb{E}\Big{[}\big{\{}\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=1)-\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=0)\mid S_{t},I_{t}=1\big{\}}f(t,S_{t})\Big{]}.$ (18)
Proof of Theorem 4.
@@
Plugging Eq. 5 from Theorem 1 into Eq. 20 and we immediately get Eq. 18. This completes the proof. ∎
@@
We restate necessary definition. As defined in Eq. 9,
$\xi(\beta):=\sum_{t=1}^{T}\omega(t)\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-\frac{p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (21)
Motivated by Robins (1999), we subtract from $\xi(\beta)$ its projection on the score functions of the treatment selection probabilities to obtain a more efficient estimating function:
@@
It follows from Lemma 6, which we will establish below, that Eq. 24 equals $\phi(\beta,\mu)=\sum_{t=1}^{T}\phi_{t}(\beta,\mu_{t})$, the proposed estimating function with improved efficiency. The terms in Eq. 25 cannot be analytically derived without imposing additional models on the relationship between current and lagged variables in the longitudinal trajectory, thus we omit them when deriving the improved estimating function. This same heuristic was employed in Cheng et al. (2023) and Bao et al. (2024). Therefore, it suffices to establish the following lemmas.
Lemma 5.
@@
which takes values $\frac{1}{p_{t}(1\mid H_{t})}$, $-\frac{1}{1-p_{t}(1\mid H_{t})}$, and 0 when $I_{t}=A_{t}=1$, $I_{t}=1$ and $A_{t}=0$, and $I_{t}=0$ (in which case $A_{t}=0$), respectively. We can directly verify that $I_{t}\dfrac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}$ also takes these exact same values under these three scenarios, respectively. This proves (d).
@@
Proof of Lemma 6.
By definition of $\xi_{t}(\beta)$ and Lemma 5(d), we have
@@
Putting together Eqs. 28, 29 and 31, we have
@@
Appendix E Proof of Asymptotic Normality (Theorem 2)
@@
Now we prove Theorem 2.
Proof of Theorem 2.
We show that the estimating function $\phi(\beta,\mu)$ is globally robust, i.e., $\mathbb{E}\{\phi(\beta^{\star},\mu)\}=0$ for all $\mu$. Then, the asymptotic normality of $\hat{\beta}$ and $\tilde{\beta}$ follows immediately from Theorems 5.1 and 5.2 of Cheng et al. (2023).
@@
The second term in Eq. 34 is 0 for any $\mu_{t}$ due to Lemma 7 and the law of iterated expectations. For the first term in Eq. 34, we have
@@
where Eq. 35 follows from Lemma 5(d) and Eq. 36 follows from Theorem 1. Plugging Eq. 36 to the first term in Eq. 34, we have
@@
where the last equality is Eq. 19. Therefore, we have proved that $\mathbb{E}\{\phi(\beta^{\star},\mu)\}=0$ for all $\mu$. We now invoke Theorems 5.1 and 5.2 of Cheng et al. (2023) to immediately get Theorem 2. ∎

### v2

@@ A.4 Example 4 (extension): DCEE captures treatment effects on future eligibility
  • $p\rho_{0}\alpha$: the negative interaction between $A_{1}$ and $A_{2}$ not accounting for the influence of $A_{1}$ on $I_{2}$;
@@
  • $p\rho_{1}\beta_{2}$: the reduced effectiveness of $A_{2}$ on $Y$ due to $A_{1}$ negatively impacting future eligiblity $I_{2}$.
Appendix B Proof of Identification Result (Theorem 2.1)
@@
We now state and prove an equivalent version of Theorem 2.1, using slightly different notation to facilitate the proof. Let $p^{1}_{t}(a\mid H_{t})$ denote the probability mass function of $A_{t}$ conditional on $H_{t}$ under the decision rule $d_{t}^{1}$: $p^{1}_{t}(a\mid H_{t})=P_{A_{t}\sim d_{t}^{1}(H_{t})}(A_{t}=a\mid H_{t})$, i.e., $p^{1}_{t}(1\mid H_{t})=I_{t}$ and $p^{1}_{t}(0\mid H_{t})=1-I_{t}$. Similarly, let $p^{0}_{t}(a\mid H_{t})$ denote the distribution of $A_{t}$ under $d_{t}^{0}$, i.e., $p^{0}_{t}(1\mid H_{t})=0$ and $p^{0}_{t}(0\mid H_{t})=1$. The following Theorem B.2 is equivalent to Theorem 2.1: one can directly verify that $\dfrac{p_{t}^{1}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}=\dfrac{{\mathds{1}}(A_{t}=I_{t})}{P(A_{t}=I_{t}\mid H_{t})}$ and $\dfrac{p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}=\dfrac{{\mathds{1}}(A_{t}=0)}{P(A_{t}=0\mid H_{t})}$ by checking equivalence under each scenario of $(A_{t},I_{t})\in\{0,1\}^{\otimes 2}$.
Theorem B.2.
Under Assumption 1, we have
Proof of Theorem B.2.
@@
Here, Eq. 16 follows from Lemma B.1, Eq. 17 follows from Assumption 1 (i), Eq. 18 follows from Assumption 1 (iii), Eq. 19 follows from Assumption 1 (iii), Eq. 20 follows from Assumption 1 (i).
The fact that Eqs. 15, 16 and 20 are equal to each other establishes Eqs. 12 and 13.
@@ Appendix C The Weighted Average Form of $\beta^{\star}$
$\times\sum_{t=1}^{T}\omega(t)\mathbb{E}(I_{t})\mathbb{E}\Big{[}\big{\{}\mathbb{E}(Y|H_{t},I_{t}=1,A_{t}=1)-\mathbb{E}(Y|H_{t},I_{t}=1,A_{t}=0)\mid S_{t},I_{t}=1\big{\}}f(t,S_{t})\Big{]}.$ (21)
Proof of Theorem C.1.
@@
Plugging Eq. 5 from Theorem 2.1 into Eq. 23 and we immediately get Eq. 21. This completes the proof. ∎
@@
We restate necessary definition. We rewrite Eq. 9 using the equivalence between Theorem 2.1 and Theorem B.2, to facilitate the derivation:
$\xi(\beta)=\sum_{t=1}^{T}\omega(t)\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-\frac{p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (24)
Motivated by Robins (1999b), we subtract from $\xi(\beta)$ its projection on the score functions of the treatment selection probabilities to obtain a more efficient estimating function:
@@
It follows from Lemma D.2, which we will establish below, that Eq. 27 equals $\phi(\beta,\mu)=\sum_{t=1}^{T}\phi_{t}(\beta,\mu_{t})$, the proposed estimating function with improved efficiency. The terms in Eq. 28 cannot be analytically derived without imposing additional models on the relationship between current and lagged variables in the longitudinal trajectory, thus we omit them when deriving the improved estimating function. This same heuristic was employed in Cheng et al. (2023) and Bao et al. (2024). Therefore, it suffices to establish the following lemmas.
Lemma D.1.
@@
which takes values $\frac{1}{p_{t}(1\mid H_{t})}$, $\frac{1}{1-p_{t}(1\mid H_{t})}$, and 0 when $I_{t}=A_{t}=1$, $I_{t}=1$ and $A_{t}=0$, and $I_{t}=0$ (in which case $A_{t}=0$), respectively. We can directly verify that $I_{t}\dfrac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}$ also takes these exact same values under these three scenarios, respectively. This proves (d).
@@
Proof of Lemma D.2.
By definition of $\xi_{t}(\beta)$ and Lemma D.1(d), we have
@@
Putting together Eqs. 31, 32 and 34, we have
@@
Appendix E Proof of Asymptotic Normality (Theorem 3.1)
@@
Now we prove Theorem 3.1.
Proof of Theorem 3.1.
We show that the estimating function $\phi(\beta,\mu)$ is globally robust, i.e., $\mathbb{E}\{\phi(\beta^{\star},\mu)\}=0$ for all $\mu$. Then, the asymptotic normality of $\widehat{\beta}$ and $\widetilde{\beta}$ follows immediately from Theorems 5.1 and 5.2 of Cheng et al. (2023).
@@
The second term in Eq. 37 is 0 for any $\mu_{t}$ due to Lemma E.1 and the law of iterated expectations. For the first term in Eq. 37, we have
@@
where Eq. 38 follows from Lemma D.1(d) and Eq. 39 follows from Theorem 2.1. Plugging Eq. 39 to the first term in Eq. 37, we have
@@
where the last equality is Eq. 22. Therefore, we have proved that $\mathbb{E}\{\phi(\beta^{\star},\mu)\}=0$ for all $\mu$. We now invoke Theorems 5.1 and 5.2 of Cheng et al. (2023) to immediately get Theorem 3.1. ∎
@@ Appendix F Numerical Computation of True Parameter Values in Simulation Studies
Appendix G Practical Considerations for Using CEE and DCEE to Inform Intervention Design
From an applied perspective, both the causal excursion effect (CEE) and the distal causal excursion effect (DCEE) offer valuable but distinct information for optimizing future digital interventions. We summarize our preliminary view of their roles and how they can be used together to inform design and evaluation below. We thank an anonymous reviewer whose question motivated this section.
G.1 Batch (offline) optimization perspective
Informing warm-start policies. Insights from CEE and DCEE could inform “warm-start” policies for future trials or deployments, potentially balancing short-term gains with long-term sustainability. For example, in a physical activity study, distal outcomes (e.g., long-term activity level) may not be the only goal; proximal outcomes (e.g., frequent activity bouts) can also improve participants’ attitudes towards activity or mental health in general. In such settings, a single scalar distal outcome may not be the only objective, and it becomes even more meaningful to consider both the distal outcome and the proximal outcomes. Designing such composite warm-start policies remains an open question, with some emerging literature in the reinforcement learning community (Shalev-Shwartz et al. 2016, Yang et al. 2024, Wu et al. 2024).
Validating proximal outcomes as surrogates. CEE focuses on short-term (proximal) responses, while DCEE captures long-term (distal) impacts. Together, they can help evaluate whether a chosen proximal outcome serves as a meaningful surrogate for the ultimate distal goal. If a strong proximal effect (CEE) does not correspond to a strong distal effect (DCEE), this suggests misalignment and motivates revision of the proximal outcome definition. Further methodology development may borrow ideas from the surrogate endpoint literature, such as focusing on proportion of treatment explained (Freedman et al. 1992), mediation-based approaches (Elliott 2023), or principal stratification-based approaches (Li et al. 2010). One challenge is the presence of time-varying treatments and the fact that the candidate “surrogacy” is a time-varying proximal outcome in MRT.
G.2 Online (real-time) optimization perspective
Reward shaping and learning speed. In real-time adaptive settings (e.g., reinforcement learning-based intervention adaptation and delivery), focusing on CEE or carefully designed proximal rewards can accelerate learning. This acts as a form of reward shaping (Ng et al. 1999, Hu et al. 2020), shifting the optimization focus from distal feedback to immediate signals, when the distal feedback is delayed.
Combining proximal and distal information. An open problem is how to best integrate CEE and DCEE in online learning settings, especially when distal outcomes are only partially observed during early stages (e.g., some participants enrolled early have distal outcomes observed, while others enrolled later only proximal data).
G.3 Feasibility considerations and iterative design approach
In early-phase or pilot trials, distal outcomes may be infeasible to collect due to long follow-up times. In such cases, CEE becomes the only feasible estimand, and primary analysis must rely solely on proximal outcomes.
An iterative design approach may involve first optimizing short-term responsiveness using CEE, then validating and refining interventions using DCEE to ensure long-term impact. Ideally, intervention design, proximal outcomes, and distal outcomes should be aligned; discrepancies among them may indicate the need to revise one or more elements. Precisely characterizing such discrepancies and determining approach revisions remain important open questions.
G.4 Insights for scientific understanding
Beyond optimizing interventions, CEE and DCEE help advance behavioral science and validate theoretical mechanisms. CEE alone indicates whether the immediate impact is as intended. DCEE alone indicates whether there is a desired long-term impact. Together, they provide evidence toward understanding potential mediation pathways, with CEE representing immediate behavioral change and DCEE capturing sustained effects. While formal mediation analysis is required for definitive causal pathway evaluation (e.g., as we preliminarily explored in Qian (2025)), jointly examining CEE and DCEE offers valuable insights.
G.5 Statistical perspective and bias-variance trade-offs
CEE and DCEE can both be viewed as parsimonious parameterizations of a fully conditional long-term effect (e.g., modeling $\mathbb{E}\{Y(\bar{a}_{T})\}$ or even $\mathbb{E}\{Y(\bar{a}_{T})\mid H_{T}(\bar{a}_{T-1})\}$ for all possible $\bar{a}_{T}$). Proximal outcomes could be easier to influence and may carry stronger signals (larger CEE), while distal outcomes could be harder to shift and may be noisier (smaller DCEE). However, one may argue that DCEE might be mroe similar to the fully conditional long-term effect given that both focus on the same distal outcome. This reflects a bias-variance trade-off spectrum:
CEEDCEEFull conditionallong-term effect Bias decreases $\rightarrow$ Variance increases $\rightarrow$ (Smallest variance, largest bias)(Largest variance, smallest bias)